please help me solve this: 1/2 + 4m = 3m - 5/2 And if possible show the check.
Answer by hyh_123(8)

Hi there.This is a very easy question and requires alot of attention.
1/2+4m=-3m-5/2
4m-3m =-1/2-5/2 (add the fractions)
m =-6/2
now simplify fraction
so you will have -3
CHECK
1/2+4(-3)=3(-3)-5/2
1/2-12 =-9-5/2
1/2+5/2 =12-9
6/2 =3
SIMPLIFY
3=3